# Developmentally-plausible Working Memory Shapes a Critical Period for Language Acquisition

### v2

@@ 5.3 Influence of Input Stimulus Length
Table 5: Accuracy in Zorro when the length of the sentence is changed
@@
The results in Table 5 reveal notable differences in model performance. For shorter sentences in the [5,10] range, NoLimit achieves slightly higher accuracy compared to DynamicLimit-Exp. However, in the [11,50] range, DynamicLimit-Exp significantly outperforms NoLimit, achieving 58.7 compared to 47.0. This suggests that DynamicLimit-Exp excels at handling moderately long sentences, likely due to its ability to dynamically adjust working memory. For longer sentences in the [51,100] and [101,150] ranges, DynamicLimit-Exp consistently outperforms NoLimit.
These findings highlight the benefits of dynamic working memory expansion in facilitating rule generalization and contextual adaptation across diverse sentence lengths. While NoLimit exhibits competitive performance on short sentences, its stagnation on longer sentences underscores its limited ability to generalize complex patterns. Conversely, DynamicLimit-Exp’s consistent performance across varying sentence lengths supports its suitability for grammatical items requiring the processing of both short and long contexts.
@@ 6 Conclusion
This study proposed a method for integrating the developmental trajectory of human working memory into the training process of LMs, inspired by the Less-is-More hypothesis. The proposed method, DynamicLimit-Exp, initially restricts working memory and gradually relaxes it exponentially during training. Experiments on both AO-CHILDES and Wikipedia showed that DynamicLimit-Exp improves grammatical learning efficiency compared to conventional methods without memory constraints or with static memory constraints. These findings suggest not only provide new approaches for developing data-efficient LMs but also offer indirect evidence supporting the CPH in human language acquisition.
@@ Acknowledgments
We are grateful to Akiyo Fukatsu for her valuable comments and suggestions. This work was supported by JSPS KAKENHI Grant Number 24H00087, JST PRESTO Grant Number JPMJPR21C2.
@@ Appendix A Details of the Training Configuration for the Base Models
Appendix B Details of Grammatical Items in Zorro
@@
Appendix C Analysis of Distributional Changes in t-SNE Space Across Training Epochs
@@ C.1 Entropy Calculation
C.2 Mean Distance Between Epochs
@@
Appendix D Development of Feature Extraction Capabilities in CASE
@@
Regarding entropy, NoLimit shows a slight decrease over time, reflecting reduced distribution diversity as training progresses. In contrast, DynamicLimit-Exp maintains or slightly increases entropy, suggesting a balanced emphasis on both basic patterns and diverse features, even in later training stages. For mean Euclidean distances between clusters, NoLimit exhibits large distances between Epoch 1 and Epoch 5 but demonstrates minimal evolution between Epoch 5 and Epoch 10. This stagnation may highlight the model’s failure to effectively generalize new rules. DynamicLimit-Exp, on the other hand, maintains substantial distances across epochs, indicating continuous embedding evolution and refinement throughout training.

### v3

@@ 5.4 Influence of Input Stimulus Length
Table 5: Accuracy in Zorro when the length of the sentence is changed. indicates a statistically significant differences (p < 0.05).
@@
The results in Table 5 reveal notable differences in model performance. For shorter sentences in the [5,10] range, NoLimit achieves slightly higher accuracy compared to DynamicLimit-Exp. However, in the [11,50] range, DynamicLimit-Exp significantly outperforms NoLimit, achieving 58.7 compared to 47.0. This suggests that DynamicLimit-Exp excels at handling moderately long sentences, likely due to its ability to dynamically adjust working memory. For longer sentences in the [51,100] and [101,150] ranges, DynamicLimit-Exp consistently outperforms NoLimit. These findings highlight the benefits of dynamic working memory expansion in facilitating rule generalization and contextual adaptation across diverse sentence lengths. While NoLimit exhibits competitive performance on short sentences, its stagnation on longer sentences underscores its limited ability to generalize complex patterns. Conversely, DynamicLimit-Exp’s consistent performance across varying sentence lengths supports its suitability for grammatical items requiring the processing of both short and long contexts.
@@ 6 Conclusion
This study proposed a method for integrating the developmental trajectory of human working memory into the training process of LMs, inspired by the Less-is-More hypothesis. The proposed method, DynamicLimit-Exp, initially restricts working memory and gradually relaxes it exponentially during training. Experiments on both AO-CHILDES and Wikipedia showed that DynamicLimit-Exp improves grammatical learning efficiency compared to conventional methods without memory constraints or with static memory constraints. These findings suggest a promising direction for building more data-efficient LMs by leveraging cognitively inspired inductive biases.
Beyond its engineering implications, this study also offers insight into the cognitive mechanisms underlying $\mathrm{L}_{1}$ acquisition. While our results do not directly demonstrate that working memory development is necessary for human learners, they serve as a computational-level plausibility test Marr (1982), showing that the hypothesized link between cognitive constraints and rule learning, central to the Less-is-More hypothesis, can be instantiated in artificial learners. Combining the observed learning efficiency gains and the cognitive plausibility of our approach, we support the hypothesis-generating idea that such developmental constraints may plausibly aid human language acquisition as well.
@@ Acknowledgments
We thank the anonymous reviewers for their helpful comments and suggestions. We are also grateful to Akiyo Fukatsu for her insightful comments and the Computational Models of Language (CoLa) Reading Group at UC Irvine for their constructive feedback. This work was supported by JSPS KAKENHI Grant Number 24H00087, JST PRESTO Grant Number JPMJPR21C2, Grant-in-Aid for JSPS Fellows JP24KJ0800, JST SPRING Grant Number JPMJSP2108.
@@ Appendix A Details of the Training Configuration for the Base Models
Appendix B Distribution of the Datasets
(a) AO-CHILDES
(b) Wikipedia
Figure 5: Word distribution of the AO-CHILDES and Wikipedia datasets used in the experiment
Figure 5 shows the sentence length distribution for the AO-CHILDES and Wikipedia datasets used in this experiment. As can be seen from the figure, AO-CHILDES, by its nature, contains more short sentences than Wikipedia.
@@ Appendix B Distribution of the Datasets
Appendix C Details of Grammatical Items in Zorro
@@
Appendix D Analysis of Distributional Changes in t-SNE Space Across Training Epochs
@@ D.1 Entropy Calculation
D.2 Mean Distance Between Epochs
@@
Appendix E Development of Feature Extraction Capabilities in CASE
@@
Table 7 shows the embedded space analysis. Regarding entropy, NoLimit shows a slight decrease over time (Epoch 1 $\xrightarrow{}$ Epoch 5), reflecting reduced distribution diversity as training progresses. In contrast, DynamicLimit-Exp maintains or slightly increases entropy, suggesting a balanced emphasis on both basic patterns and diverse features, even in later training stages. For mean Euclidean distances between clusters, NoLimit exhibits large distances between Epoch 1 and Epoch 5 but demonstrates minimal evolution between Epoch 5 and Epoch 10. This stagnation may highlight the model’s failure to effectively generalize new rules. DynamicLimit-Exp, on the other hand, maintains substantial distances across epochs, indicating continuous embedding evolution and refinement throughout training.